Help please.
Each day we purchase 160 cans of beans per second. Use the fact that there are approximatly 3.2 x 10^7 seconds in a year to approximate how many cans of beans there are purchased in one year. Write your answer in scientific notation.

Answer by jim_thompson5910(35256)

Hint: In one second, "we purchase 160 cans". In two seconds, we purchase 2*160=320 cans. In three seconds, we purchase 3*160=480 cans. This pattern continues. So in say 100 seconds, we'll purchase 100*160=16,000 cans.

This means that in seconds, we'll purchase . Ie, we're multiplying 160 and . To do this, it helps to get 160 into scientific notation.